# Learning in Compact Spaces with Approximately Normalized Transformers

### v1

@@ Abstract
In deep learning, regularization and normalization are common solutions for challenges such as overfitting, numerical instabilities, and the increasing variance in the residual stream. An alternative approach is to force all parameters and representations to lie on a hypersphere. This removes the need for regularization and increases convergence speed, but comes with additional costs. In this work, we propose a more holistic but approximate normalization (anTransformer). Our approach constrains the norm of parameters and normalizes all representations via scalar multiplications motivated by the tight concentration of the norms of high-dimensional random vectors. When applied to GPT training, we observe a 40% faster convergence compared to models with QK normalization, with less than 3% additional runtime. Deriving scaling laws for anGPT, we found our method enables training with larger batch sizes and fewer hyperparameters, while matching the favorable scaling characteristics of classic GPT architectures.
@@ 1 Introduction
Normalization techniques, such as LayerNorm, are fundamental for stable and efficient Transformer training [1, 2, 3, 4]. Loshchilov et al. [5] extends the concept and proposes the normalized Transformer (nGPT) where all latent residual representations and all parameters in the direction of the residual are normalized to lie on a hypersphere. We argue that the benefits of normalization come from two effects. Normalization prevents the representations on the residual stream from blowing up and requiring deeper layers to significantly amplify their output magnitudes. An effect observed by Sun et al. [6] and termed the “Curse of Depth” (see Figure 1). And it ensures a consistent input scale, which allows for the selection of a more suitable global learning rate.
@@
$\nu\approx(\|\bm{x}\|_{2})^{-1}\quad\text{ such that }\quad\|\nu\cdot\bm{x}\|_{2}\approx 1.$
These normalizing factors are enabled by the concentration of measure phenomenon in high-dimensional spaces, which may be perceived as a “blessing of dimensionality”. The augmentation of each operation with (approximate) normalizations, along with other modifications like bounding the norm of the input dimension of each linear map, forms the basis of our proposed approximately normalized Transformer (anTransformer). When applied to pretraining large language models, we adapt the GPT architecture to an approximately normalized GPT (anGPT) architecture without the need for additional normalization layers. Our approach does not require weight decay nor learning rate warm-up, effectively reducing the number of hyperparameters in training. When compared to vanilla GPT architecture, anGPT achieves up to $40\%$ convergence speedup. We measure a less than $3\%$ larger training step runtime, while expecting further reduction for inference times by subsuming normalization factors into the model parameters.
@@
  • We propose the approximately normalized transformer (anTranformer) in Section 4, which displays faster convergence and fewer hyperparameters at a minor increase in training time.
  • We perform extensive experimental evaluations in Section 5:
@@
      We derive hyperparameter scaling trends across multiple model sizes.
@@
      We show over $40\%$ convergence speedup compared to GPT models with QK normalization and outperform or perform on par with nGPT.
@@
      We derive compute dependent scaling laws and find that the scaling behavior matches GPT.
@@ 2 Related Work
Normalization techniques for deep neural networks have evolved significantly, from BatchNorm [7], which standardizes across batch dimensions but struggled with sequential data, to LayerNorm [1], which computes statistics across feature dimensions independently for each sample, making it ideal for transformers. The original Transformer [2] uses a Post-Norm configuration (LayerNorm after residual connections), requiring careful learning rate warm-up to prevent gradient explosion, while Pre-Norm architectures (normalization before operations) offer more stable training dynamics with higher potential learning rates [3, 4]. RMSNorm [8] further refined LayerNorm by eliminating the mean-centering step, delivering runtime improvements while maintaining performance in state-of-the-art LLMs [9, 10].
The normalized Transformer (nGPT) uses the normalized representation and parameter space to ensure that input tokens "travel" on the surface of a hypersphere, with each layer contributing a displacement towards target output predictions. These modifications render weight decay and learning rate warm-up unnecessary as vector magnitudes are explicitly controlled [5]. Empirically, nGPT demonstrates convergence speedup, reducing the number of training steps required to achieve equivalent accuracy, but comes with computational overhead due to an additional normalization layer. Further related work is given in Appendix A.
@@ 3 Preliminaries
Normalized representations are known to stabilize training and accelerate convergence [3, 4, 5]. We hypothesize that this stems primarily from two effects. First, by ensuring each layer receives a normalized input, the input scale becomes consistent across the network. This uniformity can streamline learning and simplify the selection of a global learning rate. Second, as contributions to the residual stream accumulate in vanilla Transformer architectures, subsequent layers need to amplify their output to remain influential. This leads to a growing norm of the representation on the residual stream (see Figure 1), which may destabilize training. We elaborate on both effects in the following using toy examples.
@@ 3.1 Why does normalization influence optimization?
Each entry $x_{i}$ of $\bm{x}$ converges with a rate determined by the contraction factor $\rho_{i}(\alpha)=|1-\alpha\lambda_{i}|$. To ensure convergence, we require $\rho_{i}(\alpha)<1$ with smaller values indicating faster convergence. Hence, we want set $\alpha$ to minimize the maximum contraction factor $\max_{i}\rho_{i}(\alpha)$. The $\max_{i}\rho_{i}(\alpha)$ is either achieved for the smallest $\lambda_{\min}$ or the largest $\lambda_{\max}$. To achieve the same convergence speed for both, we require the optimal learning rate $\alpha^{\star}$ to satisfy
@@
Now, consider a case where each column of $\bm{\Lambda}$ is normalized by multiplication of a diagonal preconditioner $\bm{P}=\operatorname{diag}(p_{1},\dots,p_{d})$, with $p_{j}=(\|\bm{\lambda}_{j}\|_{2})^{-1}$ where $\bm{\lambda}_{j}$ is the $j$-th column of $\bm{\Lambda}$. In this case, all $\lambda_{i}=1$ and a learning rate can be picked that leads to the fastest convergence for all coordinates. Therefore, normalization can effectively serve as a (diagonal) preconditioner and improve the convergence speed of gradient-based learning methods by making the best learning rate more effective. While the example problem is simple, it may still provide some intuition about how normalization can help learning, namely, by allowing the selection of a well-working learning rate for all parameters. Even together with methods like Adam [11], normalization may provide benefits by improving the conditioning of the optimization landscape and yielding more stable gradient statistics for moment estimation. This can be seen by the gradual increase of the variance of the first moment, see Figure B.1.
@@ 3.2 Why does the norm of the residual connection increase?
Assume independent random vectors $\bm{h}_{l}$ with $\|\bm{h}_{l}\|_{2}^{2}=1$ (for simplicity), representing the contribution of each layer $l$ on the residual connection. For the $(L+1)$-th layer to have an effective contribution to the residual state $\bm{h}_{\leq L}=\sum_{l=1}^{L}\bm{h}_{l}$, such as the ability to overwrite it, it has to have a magnitude similar to that of $\bm{h}_{\leq L}$. Specifically,
$\mathbb{E}\big{[}\|\bm{h}_{\leq L}\|_{2}^{2}\big{]}=\mathbb{E}\bigg{[}\sum_{l}^{L}\|\bm{h}_{l}\|_{2}^{2}\bigg{]}=\sum_{l}^{L}\mathbb{E}\big{[}\|\bm{h}_{l}|_{2}^{2}\big{]}=L.$
Hence, we expect that to effectively influence the residual state, $\bm{h}_{\leq L}$, $\bm{h}_{L+1}$ will produce and output with a norm of $\approx\sqrt{L}$. Such effects lead to growing outputs on the hidden state that can also be observed experimentally, see Figure 1. Since the weights are subject to regularization, the large output scales are likely produced by the scaling factor $\gamma$ in the input norm of each block of the transformer, see Figure C.1. Growing norms on the residual stream were also described in [6] and coined “Curse of Depth”. As a fix, they proposed to scale the LayerNorm output of layer $l$ by $1/\sqrt{l}$. Alternatively, this growth can also be addressed by employing Post-Norms and keeping the residual connection normalized as in Loshchilov et al. [5].
@@ 4 Approximately Normalized Transformer
Due to the potential benefits of normalization, it is tempting to normalize the inputs for each primitive, namely, linear maps, activation functions, and residual updates. Unfortunately, such excessive use of normalizations might significantly influence the runtime [5]. However, to still keep the benefits of consistent normalization at a lower cost, we explore an approach to replace normalization operations with cheaper, approximate computations.
@@ 4.1 Approximate Normalization
To reduce the overhead introduced by excessive normalization (in particular at inference time), we attempt to approximately normalize the representations in the architecture through input independent normalization factors $\nu$. Concretely, if for some vector $\bm{x}$, there exists some $\nu$, with
$\nu\approx(\|\bm{x}\|_{2})^{-1}\quad\text{ such that }\quad\|\nu\cdot\bm{x}\|_{2}\approx 1,$
we may use $\nu$ in place of $\|\bm{x}\|_{2}$ for normalizing $\bm{x}$. If we can find such normalizing factors $\nu$ for all operation primitives, we should be able to achieve approximately normalized representations without the need for exact normalization operations. It is clear that for such normalizing factors to exist, the norms of $\bm{x}$ have to concentrate closely around some value $\nu^{-1}$. Fortunately, under certain conditions, such behavior can indeed be observed.
@@
In our setting, we consider functions of the form $f(\bm{x})=\|g(\bm{x})\|_{2}$, where $g(\bm{x})$ denotes the output of a network component (e.g., a feedforward layer) given input $\bm{x}$. Assuming $g(\cdot)$ is Lipschitz and inputs $\bm{x}$ are normalized and approximately uniformly distributed on the sphere (e.g., by sampling from a Gaussian and renormalizing [12, p. 52]), the conditions of Theorem 1 are approximately satisfied. Consequently, the output norm $\|g(\bm{x})\|_{2}$ concentrates around its expected value.
While these assumptions may not hold during training, we may still observe concentration empirically, particularly due to the high dimensionality of representations. This “blessing of dimensionality” can justify approximating $\|g(\bm{x})\|_{2}$ by its expected value $\nu^{-1}:=\mathbb{E}[\|g(\bm{x})\|_{2}]$ for the purpose of normalization.
4.2 Derivation of the normalizing factors
Motivated by concentration effects in high dimensions, we derive normalizing factors for squared norms $\sqrt{\mathbb{E}[\|\bm{x}\|_{2}^{2}}]$, and then use $\nu^{-1}=\sqrt{\mathbb{E}[\|\bm{x}\|_{2}^{2}}]$. Computing $\nu$ this way should generally lead to an overestimation due to Jensen’s inequality, $\sqrt{\mathbb{E}[\|\bm{x}\|_{2}^{2}}]\geq\mathbb{E}[\sqrt{\|\bm{x}\|_{2}^{2}}]=\mathbb{E}[\|\bm{x}\|_{2}]$. However, for sufficiently high-dimensional $\bm{x}$, the bound tends to an equality due to the concentration effect. We describe the derivation for the network components used for anTransformer below. A normalization of the attention matrix is described in Appendix D since it is not part of the architecture.
@@ 4.2 Derivation of the normalizing factors
Assume $\bm{x},\bm{h}\in\mathbb{R}^{d}$ with independent entries $x_{i},h_{i}\sim\mathcal{N}(0,1)$ and normalized afterwards such that $\|\bm{x}\|_{2}=1$ and $\|\bm{h}\|_{2}=1$ We get for the classic residual update
@@
Loshchilov et al. [5] proposed to replace the classic residual update by a linear interpolation (LERP) $\bm{\bm{h}}\leftarrow\bm{h}+\bm{x}$ with $\bm{h}\leftarrow\bm{h}+\bm{\alpha}_{a}(\bm{x}-\bm{h})$ and $\alpha>0$ which leads to the benefit of explicitly learning the impact of a layer.
@@
$\bm{z}\leftarrow\bm{z}\cdot{\color[rgb]{0.0234375,0.5546875,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.0234375,0.5546875,0.171875}\bm{s_{z}}}\cdot{\color[rgb]{0.046875,0.15625,0.85546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.046875,0.15625,0.85546875}\sqrt{d}}$
@@
Table 1: Comparison between GPT implementation with SwiGLU, RMSnorm, and QK-norm (GPT+), the normalized Transformer (nGPT), and our approximated normalized GPT (anGPT). We define $\operatorname{rms}(\mathbf{h})=\mathbf{h}/\sqrt{1/n\sum_{i}^{n}h^{2}_{i}}$ and $\operatorname{norm}(\mathbf{h})=\mathbf{h}/||\mathbf{h}||_{2}$ and colored learnable parameters green, constant scaling factors blue, and normalization factors purple.
@@
As mentioned in the derivation of the linear map, we require our weights to be normalized along the input dimension. However, this does not allow to express zero vectors (the zero vector does not lie on the hypersphere). We thus only employ a bound $\|\bm{w}\|_{2}\leq 1$ similar to [13], which allows the weights to adapt more freely. Consequently, our representations are not necessarily normalized but bounded, and tending towards the boundary, see Figure E.1. We therefore refer to our representations as compact. Bounding the parameters also eliminates the need for regularization, such as weight decay. In line with Loshchilov et al. [5], we found in preliminary experiments that we do not require learning rate warm-up when initializing the parameters already normalized along the input dimension. This may be due to the more stable behavior of optimization, see e.g., Figure B.1 for the variance of the first moment in Adam.

### v2

@@ Abstract
The successful training of deep neural networks requires addressing challenges such as overfitting, numerical instabilities leading to divergence, and increasing variance in the residual stream. A common solution is to apply regularization and normalization techniques that usually require tuning additional hyperparameters. An alternative is to force all parameters and representations to lie on a hypersphere. This removes the need for regularization and increases convergence speed, but comes with additional costs. In this work, we propose a more holistic, approximate normalization via simple scalar multiplications motivated by the tight concentration of the norms of high-dimensional random vectors. Additionally, instead of applying strict normalization for the parameters, we constrain their norms. These modifications remove the need for weight decay and learning rate warm-up as well, but do not increase the total number of normalization layers. Our experiments with transformer architectures show up to 40% faster convergence compared to GPT models with QK normalization, with only 3% additional runtime cost. When deriving scaling laws, we found that our method enables training with larger batch sizes while preserving the favorable scaling characteristics of classic GPT architectures.
@@ 1 Introduction
Normalization techniques, such as LayerNorm, are fundamental for stable and efficient Transformer training [ba2016layernormalization, vaswani-neurips17a, nguyen-salazar-2019-transformers, pmlr-v119-xiong20b]. loshchilov2025ngpt extends the concept and proposes the normalized Transformer (nGPT), where all latent residual representations and all parameters in the direction of the residual are normalized to lie on a hypersphere. We argue that the benefits of normalization come from two effects. Normalization prevents the representations on the residual stream from blowing up and requiring deeper layers to significantly amplify their output magnitudes. An effect observed by sun2025curse and termed the “Curse of Depth” (see Figure 1). Additionally, normalization ensures a consistent input scale, which allows for the selection of a more suitable (global) learning rate.
@@
$\nu\approx(\|\bm{x}\|_{2})^{-1}\quad\text{ so that }\quad\|\nu\cdot\bm{x}\|_{2}\approx 1.$
These normalizing factors are enabled by the concentration of measure phenomenon in high-dimensional spaces, which may be perceived as a “blessing of dimensionality”. The augmentation of each operation with (approximate) normalizations, along with other modifications like bounding the norm of the input dimension of each linear map, forms the basis of our proposed approximately normalized Transformer (anTransformer). When applied to pretraining large language models, we adapt the GPT architecture to an approximately normalized GPT (anGPT) architecture without the need for additional normalization layers. Our approach does not require weight decay nor learning rate warm-up, effectively reducing the number of hyperparameters in training. Compared to a vanilla GPT architecture, anGPT achieves up to $40\%$ convergence speedup. Measurements show less than $3\%$ larger training step runtime, while expecting further reduction for inference times by subsuming normalization factors into the model parameters.
@@
  • The proposed approximately normalized transformer (anTransformer) in Section 4 displays faster convergence and fewer hyperparameters at a minor increase in training time.
  • Section 5 presents extensive experimental evaluations:
@@
      Hyperparameter scaling trends are derived across multiple model sizes.
@@
      Results demonstrate over 40% convergence speedup compared to GPT models with QK normalization and outperform or perform on par with nGPT.
@@
      Compute-dependent scaling laws reveal scaling behavior matching GPT.
@@ 2 Related Work
Normalization techniques for deep neural networks have evolved significantly, from BatchNorm [ioffe-icml15a], which standardizes across batch dimensions but struggled with sequential data, to LayerNorm [ba2016layernormalization], which computes statistics across feature dimensions independently for each sample, making it ideal for transformers. The original Transformer [vaswani-neurips17a] uses a Post-Norm configuration (LayerNorm after residual connections), requiring careful learning rate warm-up to prevent gradient explosion, while Pre-Norm architectures (normalization before operations) offer more stable training dynamics with higher potential learning rates [nguyen-salazar-2019-transformers, pmlr-v119-xiong20b]. RMSNorm [zhang2019rms] further refined LayerNorm by eliminating the mean-centering step, delivering runtime improvements while maintaining performance in state-of-the-art LLMs [touvron2023llama, deepseekai2024v1]. The normalized Transformer (nGPT) uses the normalized representation and parameter space to ensure that input tokens "travel" on the surface of a hypersphere, with each layer contributing a displacement towards target output predictions. These modifications render weight decay and learning rate warm-up unnecessary as vector magnitudes are explicitly controlled [loshchilov2025ngpt]. Empirically, nGPT demonstrates convergence speedups and reduces the number of training steps required to achieve equivalent accuracy. However, these improvements come with computational overhead due to additional normalization layers. Further related work is discussed in Appendix A.
@@ 3 Preliminaries
Normalized representations are known to stabilize training and accelerate convergence [nguyen-salazar-2019-transformers, pmlr-v119-xiong20b, loshchilov2025ngpt]. We hypothesize that this stems primarily from two effects. First, by ensuring each layer receives a normalized input, the input scale becomes consistent across the network. This uniformity can streamline learning and simplify the selection of a global learning rate. Second, as contributions to the residual stream accumulate in vanilla Transformer architectures, subsequent layers need to amplify their output to remain influential. This leads to a growing norm of the representation on the residual stream (see Figure 1), which may destabilize training. Both effects are elaborated below using toy examples.
@@ 3.1 Why does normalization influence optimization?
Each entry $x_{i}$ of $\bm{x}$ converges with a rate determined by the contraction factor $\rho_{i}(\alpha)=|1-\alpha\lambda_{i}|$. To ensure convergence, we require $\rho_{i}(\alpha)<1$ with smaller values indicating faster convergence. Hence, the goal is to set $\alpha$ to minimize the maximum contraction factor $\max_{i}\rho_{i}(\alpha)$. The $\max_{i}\rho_{i}(\alpha)$ is either achieved for the smallest $\lambda_{\min}$ or the largest $\lambda_{\max}$. To achieve the same convergence speed for both, we require the optimal learning rate $\alpha^{\star}$ to satisfy
@@
Now, consider a case where each column of $\bm{\Lambda}$ is normalized by multiplication of a diagonal preconditioner $\bm{P}=\operatorname{diag}(p_{1},\dots,p_{d})$, with $p_{j}=(\|\bm{\lambda}_{j}\|_{2})^{-1}$ where $\bm{\lambda}_{j}$ is the $j$-th column of $\bm{\Lambda}$. In this case, all $\lambda_{i}=1$ and a learning rate can be picked that leads to the same convergence speed for all coordinates. This fact makes good learning rates more effective, regardless of how they are found (e.g, manual tuning, grid-search, or some sophisticated optimization). Consequently, normalization can serve as a (diagonal) preconditioner and improve the convergence speed of gradient-based learning methods.
While the example problem is simple, it may still provide some intuition about how normalization can help learning, namely, by allowing the selection of a well-working learning rate for all parameters. Even together with methods like Adam [kingma-iclr15a], normalization may provide benefits by improving the conditioning of the optimization landscape and yielding more stable gradient statistics for moment estimation. This can be seen by the gradual increase of the variance of the first moment, see Figure B.1.
@@ 3.2 Why does the norm of the residual connection increase?
Assume independent random vectors $\bm{h}_{l}$ with $\mathbb{E}[\bm{h}_{l}]=0$ and $\|\bm{h}_{l}\|_{2}^{2}=1$, representing the contribution of each layer $l$ on the residual connection. For the $(L+1)$-th layer to have an effective contribution to the residual state $\bm{h}_{\leq L}:=\sum_{l=1}^{L}\bm{h}_{l}$, such as the ability to overwrite it, it has to have a magnitude similar to that of $\bm{h}_{\leq L}$. Specifically,
$\mathbb{E}\big[\|\bm{h}_{\leq L}\|_{2}^{2}\big]=\mathbb{E}\left[\left\|\sum_{l=1}^{L}\bm{h}_{l}\right\|_{2}^{2}\right]=\sum_{l=1}^{L}\mathbb{E}[\|\bm{h}_{l}\|_{2}^{2}]+\sum_{l\neq l^{\prime}}\mathbb{E}[\bm{h}_{l}]^{\top}\mathbb{E}[\bm{h}_{l^{\prime}}]=\sum_{l}^{L}\mathbb{E}\big[\|\bm{h}_{l}|_{2}^{2}\big]=L.$
Consequently, $\|\bm{h}_{L+1}\|_{2}\approx\sqrt{L}$ is expected. Such effects lead to growing outputs on the hidden state that can also be observed experimentally, see Figure 1. Since the weights are subject to regularization, the large output scales are likely produced by the scaling factor $\gamma$ in the input norm of each block of the transformer, see Figure C.1. Growing norms on the residual stream were also described in [sun2025curse] and coined “Curse of Depth”. As a fix, they proposed to scale the LayerNorm output of layer $l$ by $1/\sqrt{l}$. Alternatively, this growth can also be addressed by employing Post-Norms and keeping the residual connection normalized as in loshchilov2025ngpt.
@@ 4 Approximately Normalized Transformer
Due to the potential benefits of normalization, it is tempting to normalize the inputs for each primitive, namely, linear maps, activation functions, and residual updates. Unfortunately, such excessive use of normalizations might significantly influence the runtime [loshchilov2025ngpt]. However, to still keep the benefits of consistent normalization at a lower cost, this work explores an approach to replace normalization operations with cheaper, approximate computations.
@@ 4.1 Approximate Normalization
To reduce the overhead introduced by excessive normalization (in particular at inference time), the proposed method attempts to approximately normalize the representations in the architecture through input independent normalization factors $\nu$. Concretely, if for some vector $\bm{x}$, there exists some $\nu$, with
$\nu\approx(\|\bm{x}\|_{2})^{-1}\quad\text{ so that }\quad\|\nu\cdot\bm{x}\|_{2}\approx 1,$
we may use $\nu$ in place of $\|\bm{x}\|_{2}$ for normalizing $\bm{x}$. If such normalizing factors $\nu$ can be found for all operation primitives, approximately normalized representations should be achievable without the need for exact normalization operations. It is clear that for such normalizing factors to exist, the norms of $\bm{x}$ have to concentrate closely around some value $\nu^{-1}$. Fortunately, under certain conditions, such behavior can indeed be observed.
@@
In our setting, we consider functions of the form $f(\bm{x})=\|g(\bm{x})\|_{2}$, where $g(\bm{x})$ denotes the output of a network component (e.g., a feedforward layer) given input $\bm{x}$. Assuming $g(\cdot)$ is Lipschitz and inputs $\bm{x}$ are normalized and approximately uniformly distributed on the sphere (e.g., by sampling from a Gaussian and renormalizing [Vershynin_2018_hdpbook, p. 52]), the conditions of Theorem 1 are approximately satisfied. Consequently, the output norm $\|g(\bm{x})\|_{2}$ may concentrate around its expected value.
While these assumptions do not strictly hold during training, we may still observe concentration empirically, particularly due to the high dimensionality of representations. This “blessing of dimensionality” can justify approximating $\|g(\bm{x})\|_{2}$ by its expected value $\nu^{-1}:=\mathbb{E}[\|g(\bm{x})\|_{2}]$ for the purpose of normalization.
4.2 Derivation of the normalizing factors $\nu$
Motivated by concentration effects in high dimensions, we derive normalizing factors for squared norms $\sqrt{\mathbb{E}[\|\bm{x}\|_{2}^{2}}]$, and then use $\nu^{-1}=\sqrt{\mathbb{E}[\|\bm{x}\|_{2}^{2}}]$. Computing $\nu$ this way should generally lead to an overestimation due to Jensen’s inequality, as $\sqrt{\mathbb{E}[\|\bm{x}\|_{2}^{2}}]\geq\mathbb{E}[\sqrt{\|\bm{x}\|_{2}^{2}}]=\mathbb{E}[\|\bm{x}\|_{2}]$. However, for sufficiently high-dimensional $\bm{x}$, the bound tends to an equality due to the concentration effect. The derivation for the network components used for anTransformer is described below. A normalization of the attention matrix is described in Appendix D since it is not part of the architecture.
@@ 4.2 Derivation of the normalizing factors $\nu$
Assume $\bm{x},\bm{h}\in\mathbb{R}^{d}$ with independent entries $x_{i},h_{i}\sim\mathcal{N}(0,1)$ and normalized afterwards such that $\|\bm{x}\|_{2}=1$ and $\|\bm{h}\|_{2}=1$ . For the classic residual update, this yields
@@
loshchilov2025ngpt proposed to replace the classic residual update by a linear interpolation (LERP) $\bm{\bm{h}}\leftarrow\bm{h}+\bm{x}$ with $\bm{h}\leftarrow\bm{h}+\bm{\alpha}_{a}(\bm{x}-\bm{h})$ and $\alpha>0$ which leads to the benefit of explicitly learning the impact of a layer.
@@
[…]
@@
[…]
@@
As mentioned in the derivation of the linear map, we require our weights to be normalized along the input dimension. However, this does not allow to express zero vectors (the zero vector does not lie on the hypersphere). We thus only employ a bound $\|\bm{w}\|_{2}\leq 1$ similar to [franke2024improving], which allows the weights to adapt more freely. Consequently, our representations are not necessarily normalized but bounded, and tending towards the boundary, see Figure P.1. These representations are therefore termed compact. Bounding the parameters also eliminates the need for regularization, such as weight decay. In line with loshchilov2025ngpt, we found in preliminary experiments that we do not require learning rate warm-up when initializing the parameters already normalized along the input dimension. This may be due to the more stable behavior of optimization, see e.g., Figure B.1 for the variance of the first moment in Adam.